OpenStudy (anonymous):

The altitude of a plane as it begins to descend is shown by the linear equation a = 22,000 - 800m, where a is the plane’s altitude, in feet, and m is minutes. What is the plane’s altitude when m = 12? I think a= -9578 , but I'm not sure

OpenStudy (anonymous):

well, 12 x -800 = 9600 so a = 22000 - 9600 = 12400 your right to be not sure, a = -9578 would mean the plane is 10km underground! Well done checking :)

OpenStudy (anonymous):

(first line should be -9600 not 9600)

OpenStudy (anonymous):

So the answer is 12400 ? whew glad i checked

OpenStudy (anonymous):

yeah, what did you do to get -9578?

OpenStudy (anonymous):

I had multiplied 22,000 – 800*12 since the m = 12